birddddd:

What is the sign of the product (–7)(–2)(–5)(1)? Positive, because the products (–7)(–2) and (–5)(1) are negative, and the product of two negative numbers is positive Positive, because the products (–7)(–2) and (–5)(1) are positive, and the product of two positive numbers is positive Negative, because the product (–7)(–2) is positive, and the product (–5)(1) is negative and the product of a positive and a negative number is negative Negative, because the products (–7) (–2) and (–5) (1) are negative and the product of two negative numbers is negative

birddddd:

@ThisGirlPretty

Bearclaws72:

@shadow Can you help m8?

Bearclaws72:

Or @563blackghost

563blackghost:

\(\large\bf{- \times - = + }\) \(\large\bf{- \times + = -}\) ~~~ \(\large\bf{-7 \times -2 = 14}\) The negatives cancel to create a positive. \(\large\bf{-5 \times 1 = -5}\) A negative and a positive create a negative. \(\large\bf{14 \times -5 = -70}\) A positive and a negative create a negative.

birddddd:

k thx

birddddd:

i have mores

birddddd:

is is c

birddddd:

@563blackghost

birddddd:

is it c lol

Bearclaws72:

@Shadow

birddddd:

spam bear spam

birddddd:

oh is it c

Bearclaws72:

Shut hell up buttfeed do your work

Bearclaws72:

@Shadow Would the asnwer that Blackghost gave him be C?

birddddd:

i got c by my self

birddddd:

i have no clue what blackghost said lol

563blackghost:

yea it's c.....

birddddd:

yeah boi

Bearclaws72:

cringe^

563blackghost:

does the latex not work for you @birddddd ?

birddddd:

no

563blackghost:

oh ok, ill make sure to write it out without next time :)

birddddd:

for some reason i cant see it

birddddd:

thank you

563blackghost:

−×−=+ −×+=− ~~~ −7×−2=14 The negatives cancel to create a positive. −5×1=−5 A negative and a positive create a negative. 14×−5=−70 A positive and a negative create a negative. Can you see it now?

birddddd:

yup

birddddd:

thx

563blackghost:

np